# Graph Classification Gaussian Processes via Hodgelet Spectral Features

### v1

@@ C.3 Hodge decomposition
The Helmholtz decomposition, a famous calculus theorem often called the fundamental theorem of vector calculus, states that any 2D vector fields on a Euclidean domain can be expressed as the sum of two orthogonal components: (1) one that intuitively has no sinks or sources, called divergence-free or solenoidal, and (2) one that has no vortices, also called curl-free or irrotational. This decomposition extended to differential forms on manifolds is referred to as the Hodge decomposition. In the case of Riemannian manifolds, we can identify differential 1-forms with vector fields, which results in a decomposition of vector fields on manifolds. The Hodge decomposition implies that the $k$-simplex feature space $\mathbb{R}^{N_{k}}$ (for a single channel) can be decomposed into the sum of three orthogonal components: exact, co-exact, and harmonic. This leads us to the decomposition
@@
where $\text{im}(\boldsymbol{B}_{k+1})$ is called the exact subspace, $\text{im}(\boldsymbol{B}_{k}^{\top})$ the co-exact subspace and $\text{ker}(\boldsymbol{L}_{k})$ the harmonic subspace. Interestingly, we can see that the dimension of $\text{ker}(\boldsymbol{L}_{k})$ is equal to the $k$-th Betti number, which describe the number of $k$-dimensional holes. The Hodge decomposition implies the eigendecomposition of the Hodge Laplacian,
@@
where $(\boldsymbol{\Lambda}_{ke},\boldsymbol{U}_{ke})$ are the non-zero eigenvalues and eigenvectors of $\boldsymbol{B}_{k}^{\top}\boldsymbol{B}_{k}$, $(\boldsymbol{\Lambda}_{kc},\boldsymbol{U}_{kc})$ the non-zero eigenvalues and eigenvectors of $\boldsymbol{B}_{k+1}\boldsymbol{B}_{k+1}^{\top}$, and $(\boldsymbol{\Lambda}_{kh},\boldsymbol{U}_{kh})$ the zero eigenvalues and eigenvectors of $\boldsymbol{L}_{k}$. We observe that the exact eigenbasis $\boldsymbol{U}_{0e}=\boldsymbol{0}$, since $\boldsymbol{B}_{0}=\boldsymbol{0}$. Similarly, co-exact eigenbasis $\boldsymbol{U}_{k+1e}=\boldsymbol{0}$, since $\boldsymbol{B}_{k+1}=\boldsymbol{0}$.
@@
In the same spirit as Section 2.1, we define the wavelet transform on $k$-simplex features, obtain wavelet spectral features, and finally define the Hodgelet kernel on simplicial $K$-complexes,

### v2

@@ C.3 Hodge decomposition
The Helmholtz decomposition is often called the fundamental theorem of vector calculus. This famous theorem states that any 2D vector fields on a Euclidean domain can be expressed as the sum of two orthogonal components: (1) one that intuitively has no sinks or sources, called divergence-free or solenoidal, and (2) one that has no vortices, sometimes called curl-free or irrotational. This decomposition extended to differential forms on manifolds is referred to as the Hodge decomposition. For Riemannian manifolds, we can identify differential $1$-forms with vector fields. This implies that Hodge decomposition can be interpreted as a decomposition of vector fields on Riemannian manifolds. The Hodge decomposition on the $k$-simplex feature space $\mathbb{R}^{N_{k}}$ (for a single channel) leads to the sum of three orthogonal components: exact, co-exact, and harmonic,
@@
where $\text{im}(\boldsymbol{B}_{k+1})$ is the exact subspace, $\text{im}(\boldsymbol{B}_{k}^{\top})$ is the co-exact subspace, and $\text{ker}(\boldsymbol{L}_{k})$ is the harmonic subspace. Interestingly, we note that the dimension of $\text{ker}(\boldsymbol{L}_{k})$ is equal to the $k$-th Betti number, which describe the number of $k$-dimensional holes. The Hodge decomposition implies the eigendecomposition of the Hodge Laplacian,
@@
where $(\boldsymbol{\Lambda}_{ke},\boldsymbol{U}_{ke})$ are the non-zero eigenvalues and eigenvectors of $\boldsymbol{B}_{k}^{\top}\boldsymbol{B}_{k}$, $(\boldsymbol{\Lambda}_{kc},\boldsymbol{U}_{kc})$ are the non-zero eigenvalues and eigenvectors of $\boldsymbol{B}_{k+1}\boldsymbol{B}_{k+1}^{\top}$, and $(\boldsymbol{\Lambda}_{kh},\boldsymbol{U}_{kh})$ are the zero eigenvalues and eigenvectors of $\boldsymbol{L}_{k}$. We observe that the exact eigenbasis $\boldsymbol{U}_{0e}=\boldsymbol{0}$ as $\boldsymbol{B}_{0}=\boldsymbol{0}$. Similarly, co-exact eigenbasis $\boldsymbol{U}_{k+1e}=\boldsymbol{0}$ since $\boldsymbol{B}_{k+1}=\boldsymbol{0}$. We also have that $\boldsymbol{U}_{ke}$ spans $\text{im}(\boldsymbol{B}_{k+1})$, $\boldsymbol{U}_{kc}$ spans $\text{im}(\boldsymbol{B}_{k}^{\top})$, and $\boldsymbol{U}_{kh}$ spans $\text{ker}(\boldsymbol{L}_{k})$.
@@
In the same spirit as in Section 2.1, we compute the wavelet transform on the $k$-simplex features and then obtain the wavelet spectral features. Finally, we present the Hodgelet kernel on simplicial $K$-complexes,
@@ D.2 Vector Field Classification Details
(a) Divergence-free field
(b) Regular triangular mesh